Studies show that massage therapy has a variety of health benefits and it is not too expensive. A sample of 11 typical one-hour massage therapy sessions showed an average charge of $60. The population standard deviation for a one-hour session is σ-$5.45 a. What assumptions about the population should we be willing to make if a margin of error is desired? Select your answer b. Using 90% confidence, what is the margin of error (to 2 decimals)? c. Using 95% confidence, what is the margin of error (to 2 decimals)?
Answer #1

Answer

(A) Assumption that we need to make about the population when we want to calculate the margin of error is that "the population distribution is at least approximately normal"

This means that the population must be distributed normally in order to solve for the margin of error.

(B) we have sample size n =11 and standard deviation = 5.45

z score for 90% confidence level is 1.64 (using z distribution table)

Formula for margin of error= ME = z*(sigma/sqrt{n})

setting the given values, we get

ME = 1.64 * (5.45/V11) 1.64 * 1.643 2.69

So, required margin of error is 2.69 (rounded to 2 decimals)

(C) we have sample size n =11 and standard deviation = 5.45

z score for 95% confidence level is 1.96 (using z distribution table)

Formula for margin of error= ME = z*(sigma/sqrt{n})

setting the given values, we get

ME = 1.96 * (5.45/V11) 1.96 * 1.643 3.22

So, required margin of error is 3.22 (rounded to 2 decimals)